OpenStudy (anonymous):

Evaluate: x^3 - 2x^2 - 4x when x = -2 I've got it to 8, but apparently it's -8. Could somebody explain please.

OpenStudy (austinl):

\(x^3-2x^2-4x,~x=-2\) \((-2)^3-2(-2)^2-4(-2)=~?\) \(-8-8+8=~?\)

OpenStudy (anonymous):

Stupid mistake on my part, thanks again.

OpenStudy (austinl):

No problemo!